1)solve both by elimination method: 5x+6y=2
10x+12y=4

2) 2r-7s=-12
7r+2s=11

Found 2 solutions by solver91311, stanbon:
Answer by solver91311(24713)







Multiply the first equation by -2



Add the two equations term by term resulting in the triviality:



The system is consistent and dependent. The two equations graph to the same line.

John

My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism


Answer by stanbon(75887)

1)solve both by elimination method:
5x+6y=2
10x+12y=4
----------------
Multiply thru the 1st equation by 2 to get:
10x+12y = 4
---
Subtracting that from the 2nd equation gives:
0 = 0
---
Conclusion: All points where y = (-5/6)x +(1/3)
===============================================
2)
2r-7s=-12
7r+2s=11
============
Multiply thru 1st by 7
Multiply thru 2nd by 2
-----
14r - 49s = -84
14r + 4s = 22
--------------------
Subtract and solve for "s":
53s = -106
s = -2
---
Solve for "r":
2r - 7*(-2) = -12
2r+14 = -12
2r = -26
r = -13
===============
Cheers,
Stan H.
================